# An algebraic characterization for spectral radii of trees

### v2

@@ 1 Introduction
In order to prove Theorem 1.5, by Corollary 3.3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $[r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$. We start with the following simple result.
Lemma 4.1.
@@
    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval $[3m+t-1,4m+t]$.
@@
(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $10\leq 4m+t-s\leq m+$ 1 and
@@
Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 4.1.
@@
Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $20\leq i\leq m-$ 2.
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any $i\in{\mathbb{N}}^{+}$ with $21\leq i\leq m-$ 2, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
@@
By Lemma 4.1, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $20\leq i\leq m-$ 2, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
@@
Hence Claim 4 holds for $22m\leq s\leq 3m-$ 2. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
For any $i\in{\mathbb{N}}^{+}$ with $10\leq i\leq m-$ 1, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
@@
Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 4.1, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
@@
By Lemma 4.1 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 4.4.
@@
Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$. By Lemma 4.1, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 1 holds for $i=1$.
@@
Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $-2\leq i\leq m+1$.
@@
Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $-2\leq i\leq 1$.
@@
By Lemma 4.1 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 4.5.
@@
Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $11\leq i\leq t-$ 1. Then, by definition,
@@
Claim 2: For any integer $i$ with $-3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
@@
Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
By Lemma 4.1, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $33m+2\leq s\leq 4m+$ 3.
@@
Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $12m\leq s\leq 3m+$ 1. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
We are now going to prove Theorem 1.5.
Proof of Theorem 1.5: The necessity follows from (1). Now we prove the sufficiency.
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 4.1 and Propositions 4.24.34.4 and 4.5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds. Note that
@@
By Corollary 3.3, for any integer $s$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\lambda(T)=\sqrt{k}$. $\Box$
@@ 5 Further study
Recall that $\Lambda$ is the set of $\lambda(T)$ over all trees $T$ of order at least $2$. Conjecture 1.4 asks an interesting question on the existence of a tree $T$ with $\lambda(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\Lambda\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
Another interesting problem is on the numbers contained in $\Lambda$. Let ${\cal P}$ be the set of real-rooted monic polynomials with integer coefficients. Salez [21] showed that every totally real algebraic integer (i.e. a root of some polynomial in ${\cal P}$) is an eigenvalue of some finite tree. Now let $\Lambda^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
Problem 5.1.
Is there a positive real number $c$ such that $\{\alpha\in\Lambda^{\prime}:\alpha\geq c\}\subseteq\Lambda$?
It can be proved by applying Theorem 1.1 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\Lambda$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
Proposition 5.2.
For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
Proof. Assume that $p-q=4d$. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
@@
By Corollary 3.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$.
Problem 5.3.
Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $4\big{|}(p-q)$?
Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$. Then, (1) yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. From the proving of Lemma 2.3, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 5.4.
Let $\lambda\in\Lambda$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?

### v3

@@ 4 Proof of Theorem 3
In order to prove Theorem 3, by Corollary 3, it suffices to show that $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ ($\subseteq{\mathscr{W}}_{s}(\sqrt{k})$) for all integers $s$ in the interval $r_{0},k]$, since this implies that $s\times\left(\frac{s}{\sqrt{k}}\right)^{-1}=\sqrt{k}$. We start with the following simple result.
Lemma 4.
@@
    when $m\geq 1$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{m+2}(\sqrt{k})$ for each integer $s$ in the interval [3m+t1,4m+t][3m+t-1,4m+t].
@@
(ii) Assume that $m\geq 1$. For any $s\in{\mathbb{N}}^{+}$ with $3m+t-1\leq s\leq 4m+t$, we have $0\leq 4m+t-s\leq m+1$ and
@@
Proof. We will prove that the conclusion in the following claims. This result is true for integers $s$ in the interval $[3m,4m]$ by Lemma 4.
@@
Claim 3: $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ for all integers $i$ with $0\leq i\leq m-2$.
By Claim 2, $\frac{m-1}{m}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$. For any $i\in{\mathbb{N}}^{+}$ with $1\leq i\leq m-2$, if $\frac{m-i}{m-i+1}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, by definition,
@@
By Lemma 4, Claim 4 holds for $3m-1\leq s\leq 4m$. For any $i\in{\mathbb{N}}^{+}$ with $0\leq i\leq m-2$, we have $\frac{m-i-1}{m-i}\sqrt{m}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$ by Claim 3, and thus
@@
Hence Claim 4 holds for $2m\leq s\leq 3m-2$. In particular, $\sqrt{m}=\frac{2m}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$.
For any $i\in{\mathbb{N}}^{+}$ with $0\leq i\leq m-1$, if $\frac{2m-i}{2\sqrt{m}}\in{\mathscr{W}}_{m+2}(2\sqrt{m})$, then
@@
Let $Q(i)=\frac{(2i+1)m}{i\sqrt{4m+1}}$. By Lemma 4, $Q(1)=\frac{3m}{\sqrt{4m+1}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+1}).$ By definition, if $Q(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+1})$, then,
@@
By Lemma 4 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 4.
@@
Let $P(i)=\frac{(2i+1)m+i}{i\sqrt{4m+2}}$ and $Q(i)=\frac{(2i+1)m+i+1}{i\sqrt{4m+2}}$. By Lemma 4, $P(1)=\frac{3m+1}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ and $Q(1)=\frac{3m+2}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$. Thus, Claim 1 holds for $i=1$.
@@
Claim 2: $\frac{2m-i}{\sqrt{4m+2}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for integers $i$ with $2\leq i\leq m+1$.
@@
Thus, $W(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+2})$ for each integer $i$ with $2\leq i\leq 1$.
@@
By Lemma 4 and Claims 2 and 3, the conclusion holds. $\Box$
Proposition 5.
@@
Now assume that $t$ is an integer with $t\geq 2$ such that $f_{1}(i),f_{2}(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for all integers $i$ with $1\leq i\leq t-1$. Then, by definition,
@@
Claim 2: For any integer $i$ with $3\leq i\leq m$, $\frac{2m-i}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
@@
Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $2m\leq s\leq 4m+3$, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
By Lemma 4, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $3m+2\leq s\leq 4m+3$.
@@
Thus, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ for each integer $s$ with $2m\leq s\leq 3m+1$. Hence Claim 3 holds. The result then follows from Claims 2 and 3. $\Box$
We are now going to prove Theorem 3.
Proof of Theorem 3: The necessity follows from Theorem 1. Now we prove the sufficiency.
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 4 and Propositions 234 and 5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds. Note that
@@
By Corollary 3, for any integer $s$ in $r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\rho(T)=\sqrt{k}$. $\Box$
@@ 5 Further study
Recall that $\mathbb{P}$ is the set of $\rho(T)$ over all trees $T$ of order at least $2$. Problem 1 asks an interesting question on the existence of a tree $T$ with $\rho(T)=\alpha$ and $\Delta(T)=r$ for any given $\alpha\in\mathbb{P}\setminus\{1\}$ and integer $r$ satisfying $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$.
Another interesting problem is on the numbers contained in $\mathbb{P}$. Let $\cal P$} be the set of real-rooted monic polynomials with integer coefficients. Salez [28] showed that every totally real algebraic integer (i.e. a root of some polynomial in $\cal P$}) is an eigenvalue of some finite tree. Now let $\mathbb{P}^{\prime}$ be the set of real numbers $\alpha$ such that $\alpha$ is the largest real root of some $P\in{\cal P}$.
Problem 2.
Is there a positive real number $c$ such that $\{\alpha\in\mathbb{P}^{\prime}:\alpha\geq c\}\subseteq\mathbb{P}$?
It can be proved by applying Theorem 2 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ if 4|(pq)4|(p-q), $\sqrt{p+\sqrt{q}}\in\mathbb{P}$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\mathbb{P}$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ whenever 4|(pq)4|(p-q).
Proposition 6.
For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if pqp-q is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\rho(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
Proof. Assume that pq=4dp-q=4d. Let $r=\max\{q,d\}+1$ and $\alpha=\frac{\sqrt{p}+\sqrt{q}}{2}$. By definition,
@@
By Corollary 4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\rho(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if 4|(pq)4|(p-q) is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\mathbb{P}$.
Problem 3.
Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\mathbb{P}$ if and only if $4\big{|}(p-q)$?
Let $T$ be any tree with $\alpha=\rho(T)$ and $r=\Delta(T)$. Then, Theorem 1 yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. From the proving of Lemma 1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 4.
Let $\lambda\in\mathbb{P}$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?